Circular No. 7601

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2001 BB_{50} (LINEAR-NEAT) S. Pravdo, K. Lawrence, and E. Helin, Jet Propulsion Laboratory, re-

ported the discovery of a comet on Mar. 20 CCD images taken with the NEAT 1.2-m reflector at Haleakala, the object showing a short eastward tail, a nuclear condensation of size < 3′′, and a coma diameter of ∼ 10′′. T. B. Spahr, Minor Planet Center, linked this object first with an object reported as asteroidal by LINEAR on Mar. 18 (m_{2} = 19.5) and then to the LINEAR object 2001 BB_{50}, observed on Jan. 21 and 26 (MPS 25734). Following posting on the NEO Confirmation Page, C. Jacques, Belo Horizonte, Brazil, also reported a 10′′ coma and m_{1} = 18.6 on CCD images taken on Mar. 21 (0.3-m reflector). Full astrometry and the orbital elements given below appear on MPEC 2001-F26:

2001 UT α_{2000} δ_{2000} m_{1} Observer
Jan. 21.46537 13^{h}36^{m}56.^{s}92 −11^{o}30′13′′.4  LINEAR
Mar. 20.61113 14 03 14.84 −20 22 25.6 18.4 NEAT
T = 2001 Jan. 30.3429 TT  ω = 189.^{o}3139
e = 0.587682  Ω = 355.8217 }2000.0
q = 2.347221 AU  i = 10.6196
a = 5.692747 AU n^{o} = 0.0725640 P = 13.583 years

COMET C/2001 C6 (SOHO) Further to IAUC 7585, D. Biesecker reports his measurements for a

comet in images obtained by the C2 and C3 coronagraphs and found by M. Boschat on SOHO website images. The reduced observations and orbital elements by B. G. Marsden appear on MPEC 2001-D42.

2001 UT  α_{2000}  δ_{2000}
Feb. 13.888  22^{h}15.^{m}4  −15^{o}34′

SUPERNOVAE 2001W AND 2001aa R. Chornock, A. V. Filippenko, and M. Modjaz, University of Califor-

nia at Berkeley, report that inspection of a CCD spectrum (range 320–1000 nm) reveals that SN 2001W (IAUC 7590) is of type II. Well-developed P-Cyg profiles are seen in the hydrogen Balmer lines, and Fe II absorption is present. SN 2001aa (IAUC 7593) is also of type II, with a spectrum (range 550–1000 nm) similar to that of SN 2001W.

2001 March 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green